# JanusDNA: A Powerful Bi-directional Hybrid DNA Foundation Model

### v1

@@ 3.1 Bidirectional Efficient Training
Here, $q_{\text{idx}}$ and $kv_{\text{idx}}$ are the 0-indexed indices of the query and key-value pairs within the $2T$-length $R_{\text{input}}$, respectively. $T$ is the original sequence length. The mask $\mathcal{M}_{ij}$ is a binary matrix where allowed attentions are 1 and disallowed are 0 (or $\infty$- ∞ after softmax). The first two cases handle causal attention within the $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, respectively. The third and fourth cases manage the cross-attention between $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, precisely controlling information flow to maintain the integrity of bidirectional prediction without future peeking relative to the token being predicted. The output of this attention mechanism provides a fused representation $H^{\text{final}}_{t}$ for each token $x_{t}$, which is then used to make the final prediction, optimizing $\mathcal{L}_{\text{bidirectional}}$.
@@
To empirically assess the learning efficiency of Janus modeling against conventional bidirectional approaches – masked modeling, we conducted a comparative experiment focused on last-token prediction. This task was chosen as it allows a direct comparison: both a standard masked language model and our Janus modeling approach can predict the final token $x_{T}$ given the preceding context $x_{1},\ldots,x_{T-1}$, ensuring a fair basis for evaluation.
For this demonstration, we configured two model variants: Janus Model: A single-layer bidirectional Mamba architecture equipped with a FlexAttention layer pytorch_flexattention for bidirectional fusion, utilizing the Janus-specific attention mask $\mathcal{M}_{ij}$ (Equation 6). This model predicts $x_{t}$ based on the fused bidirectional context of $x_{1},\ldots,x_{t-1},x_{t+1},\ldots,x_{T}$ as per the Janus modeling objective. Masked Language Model: As a baseline, we construct a comparable single-layer Mamba architecture followed by the same FlexAttention layer without an attention mask, trained using the conventional masked language modeling (MLM) objective, where a fraction of tokens (typically 15%) contribute to the loss.
@@ 4.2 Downstream Tasks
We evaluate our model on three different benchmarks: Genomic Benchmark genomicbenchmark , Nucleotide Transformer Benchmark nucleotidebenchmark , and DNALONGBENCH chengdna . We follow all benchmark settings of Genomic Benchmark and Nucleotide Transformer Benchmark as described in caduceus . Accordingly, we adopt the reported results from caduceus as our reference. Considering the practical computational cost of sparse MoEs, we adjust the model’s hidden size to match or slightly reduce the number of activated parameters compared to the baseline caduceus , ensuring a fair comparison. As we introduce the model with a middle attention layer in the ablation experiments (Section C), we present results for models both with and without mid-attention on the Genomic Benchmark and Nucleotide Transformer Benchmark.
@@ 4.2.1 Genomic Benchmark
Mouse Enhancers $0.715$±$\pm 0.087$ $0.780$±$\pm 0.025$ $0.743$±$\pm 0.054$ $0.754$±$\pm 0.074$ $0.793$±$\pm 0.058$ $0.752$±$\pm 0.041$ $0.750$±$\pm 0.031$
Coding vs. Intergenomic $0.892$±$\pm 0.008$ $0.904$±$\pm 0.005$ $0.904$±$\pm 0.004$ $0.915$±$\pm 0.003$ $0.910$±$\pm 0.003$ $0.913$±$\pm 0.001$ $0.914$±$\pm 0.002$
Human vs. Worm $0.942$±$\pm 0.002$ $0.964$±$\pm 0.002$ $0.967$±$\pm 0.002$ $0.973$±$\pm 0.001$ $0.968$±$\pm 0.002$ $0.972$±$\pm 0.001$ $0.971$±$\pm 0.001$
Human Enhancers Cohn $0.702$±$\pm 0.021$ $0.729$±$\pm 0.014$ $0.732$±$\pm 0.029$ $0.747$±$\pm 0.004$ $0.745$±$\pm 0.007$ $0.745$±$\pm 0.004$ $0.740$±$\pm 0.002$
Human Enhancer Ensembl $0.744$±$\pm 0.122$ $0.849$±$\pm 0.006$ $0.862$±$\pm 0.008$ $0.893$±$\pm 0.008$ $0.900$±$\pm 0.006$ $0.899$±$\pm 0.003$ $0.900$±$\pm 0.005$
Human Regulatory $0.872$±$\pm 0.005$ $0.869$±$\pm 0.012$ $0.814$±$\pm 0.211$ $0.872$±$\pm 0.011$ $0.873$±$\pm 0.007$ $0.877$±$\pm 0.002$ $0.870$±$\pm 0.006$
Human OCR Ensembl $0.698$±$\pm 0.013$ $0.783$±$\pm 0.007$ $0.815$±$\pm 0.002$ $0.828$±$\pm 0.006$ $0.818$±$\pm 0.006$ $0.825$±$\pm 0.002$ $0.825$±$\pm 0.003$
Human NonTATA Promoters $0.861$±$\pm 0.009$ $0.944$±$\pm 0.002$ $0.933$±$\pm 0.007$ $0.946$±$\pm 0.007$ $0.945$±$\pm 0.010$ $0.949$±$\pm 0.006$ $0.945$±$\pm 0.006$
@@
Enhancer $0.451$±$\pm 0.108$ $0.516$±$\pm 0.098$ $0.548$±$\pm 0.144$ $0.517$±$\pm 0.117$ $0.546$±$\pm 0.073$ $0.564$±$\pm 0.022$ $0.535$±$\pm 0.036$
Enhancer types $0.309$±$\pm 0.134$ $0.423$±$\pm 0.051$ $0.424$±$\pm 0.132$ $0.386$±$\pm 0.185$ $0.439$±$\pm 0.054$ $0.462$±$\pm 0.049$ $0.470$±$\pm 0.025$
Promoter: All $0.954$±$\pm 0.006$ $0.971$±$\pm 0.006$ $0.976$±$\pm 0.006$ $0.960$±$\pm 0.005$ $0.970$±$\pm 0.004$ $0.969$±$\pm 0.002$ $0.971$±$\pm 0.002$
     NonTATA $0.955$±$\pm 0.010$ $0.972$±$\pm 0.005$ $0.976$±$\pm 0.005$ $0.959$±$\pm 0.008$ $0.969$±$\pm 0.011$ $0.971$±$\pm 0.003$ $0.971$±$\pm 0.002$
     TATA $0.960$±$\pm 0.023$ $0.955$±$\pm 0.021$ $0.966$±$\pm 0.013$ $0.944$±$\pm 0.040$ $0.953$±$\pm 0.016$ $0.956$±$\pm 0.010$ $0.958$±$\pm 0.008$
@@ 4.2.3 DNA Long Range Benchmark
To further assess our model’s ability to capture long-range dependencies in DNA sequences, we evaluate it on the expression Quantitative Trait Loci (eQTL) prediction task from DNALONGBENCH chengdna . The eQTL task measures whether a nucleotide variant can influence the expression of a target gene based on the sequences of the gene and its surrounding regions.
@@
Artery Tibial $0.741$ $0.690$ $0.803$
Adipose Subcutaneous $0.736$ $0.759$ $0.741$
Cells Cultured fibroblasts $0.639$ $0.690$ $0.771$
Muscle Skeletal $0.621$ $0.789$ $0.803$
Nerve Tibial $0.683$ $0.842$ $0.877$
Skin Not Sun Exposed Suprapubic $0.710$ $0.812$ $0.875$
Skin Sun Exposed Lower leg $0.700$ $0.692$ $0.706$
Thyroid $0.612$ $0.703$ $0.752$
Whole Blood $0.689$ $0.769$ $0.794$
Due to limited computational resources, we compare our model against the state-of-the-art DNA language model, Caduceus-PH caduceus , which is also trained with the same data scale for fair comparison, and the expert model for this task, Enformer enformer . The results for Enformer are taken directly from DNALONGBENCH chengdna . For Caduceus-PH, we fine-tune the official HuggingFace-released weights, which are pretrained on sequences of length 131k. Our model is pretrained and fine-tuned using the same setup and sequence length as Caduceus-PH to ensure a fair comparison.
@@ Appendix C Ablation Experiments
To determine the optimal hybrid architecture for DNA sequence modeling that effectively balances local and global attention, we perform ablation experiments on various configurations of the unidirectional encoder (i.e., modules preceding the bidirectional fusion layer). Referring to [21], we also explore the value of the additional mid-attention layer. Specifically, we evaluate the following configurations: 1) mamba and FFN blocks only, 2) mamba and FFN blocks with mid-attention, 3) mamba and FFN blocks with MoE, and 4) mamba and FFN blocks with both mid-attention and MoE. The ratio of MoE to replace FFNs is set to 0.5. In models with mid-attention, the two independent bidirectional Mamba blocks at the 4th layer are replaced with two independent causal Attention blocks implemented using FlexAttention [23].
The models are pre-trained on sequences of lengths 1024 and 131072, with batch sizes of 128 and 1, respectively, using a single GPU. All other hyperparameters and training settings are consistent with the pre-training setup described earlier. The training perplexity results for these configurations are shown in Figure 4.
@@
5, indicate that models with mid-attention perform better at a hidden dimension of 32. However, as the hidden dimension increases, models without mid-attention outperform those with mid-attention at a dimension of 72. At a dimension of 128, the performance of models with and without mid-attention becomes comparable. These findings suggest that mid-attention provides diminishing benefits as the hidden dimension grows larger, eventually becoming negligible at higher dimensions.

### v3

@@ 3.1 Bidirectional Efficient Training
Here, $q_{\text{idx}}$ and $kv_{\text{idx}}$ are the 0-indexed indices of the query and key-value pairs within the $2T$-length $R_{\text{input}}$, respectively. $T$ is the original sequence length. The mask $\mathcal{M}_{ij}$ is a binary matrix where allowed attentions are 1 and disallowed are 0 (or $\infty$- ∞ after softmax). The first two cases handle causal attention within the $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, respectively. The third and fourth cases manage the cross-attention between $\mathcal{R}_{\text{fwd}}$ and $\mathcal{R}_{\text{bwd}}$ segments, precisely controlling information flow to maintain the integrity of bidirectional prediction without future peeking relative to the token being predicted. The output of this attention mechanism provides a fused representation $H^{\text{final}}_{t}$ for each token $x_{t}$, which is then used to make the final prediction following a repositioning step, where the representations of the same token are summed, except for the first and last tokens, due to their representations containing information from only one direction, while optimizing $\mathcal{L}_{\text{bidirectional}}$.
@@
To empirically assess the learning efficiency of Janus modeling against conventional bidirectional approaches, masked modeling, we conducted a comparative experiment focused on last-token prediction. This task was chosen as it allows a direct comparison: both a standard masked language model and our Janus modeling approach can predict the final token $x_{T}$ given the preceding context $x_{1},\ldots,x_{T-1}$, ensuring a fair basis for evaluation.
Figure 3: Janus and Masked Modeling Efficiency Validation. Both models are pre-trained from scratch using identical hyperparameter settings, with the only difference being the masking strategy applied in the final fusion attention layer. Last-token prediction is used to enable a fair comparison of learning efficiency between the two models.
@@
For this demonstration, we configured two model variants as Figure 3: Janus Model: A single-layer bidirectional Mamba architecture equipped with a FlexAttention layer pytorch_flexattention for bidirectional fusion, utilizing the Janus-specific attention mask $\mathcal{M}_{ij}$ (Equation 6). This model predicts $x_{t}$ based on the fused bidirectional context of $x_{1},\ldots,x_{t-1},x_{t+1},\ldots,x_{T}$ as per the Janus modeling objective. Masked Language Model: As a baseline, we construct a comparable single-layer Mamba architecture followed by the same FlexAttention layer without an attention mask, trained using the conventional masked language modeling (MLM) objective, where a fraction of tokens (typically 15%) contribute to the loss.
@@ 4.2 Downstream Tasks
We evaluate our model on three different benchmarks: Genomic Benchmark genomicbenchmark , Nucleotide Transformer Benchmark nucleotidebenchmark , and DNALONGBENCH cheng2025dnalongbench . We follow all benchmark settings of Genomic Benchmark and Nucleotide Transformer Benchmark as described in caduceus . Accordingly, we adopt the reported results from caduceus as our reference. Considering the practical computational cost of sparse MoEs, we adjust the model’s hidden size to match or slightly reduce the number of activated parameters compared to the baseline caduceus , ensuring a fair comparison. As we introduce the model with a middle attention layer in the ablation experiments (Section C), we present results for models both with and without mid-attention on the Genomic Benchmark and Nucleotide Transformer Benchmark.
@@ 4.2.1 Genomic Benchmark
Mouse Enhancers $0.715$±$\pm 0.087$ $0.780$±$\pm 0.025$ $0.743$±$\pm 0.054$ $0.754$±$\pm 0.074$ $\bf\bf 0.793$±$\pm 0.058$ $0.752$±$\pm 0.041$ $0.750$±$\pm 0.031$
Coding vs. Intergenomic $0.892$±$\pm 0.008$ $0.904$±$\pm 0.005$ $0.904$±$\pm 0.004$ $\bf 0.915$±$\pm 0.003$ $0.910$±$\pm 0.003$ $0.913$±$\pm 0.001$ $0.914$±$\pm 0.002$
Human vs. Worm $0.942$±$\pm 0.002$ $0.964$±$\pm 0.002$ $0.967$±$\pm 0.002$ $\bf 0.973$±$\pm 0.001$ $0.968$±$\pm 0.002$ $0.972$±$\pm 0.001$ $0.971$±$\pm 0.001$
Human Enhancers Cohn $0.702$±$\pm 0.021$ $0.729$±$\pm 0.014$ $0.732$±$\pm 0.029$ $\bf 0.747$±$\pm 0.004$ $0.745$±$\pm 0.007$ $0.745$±$\pm 0.004$ $0.740$±$\pm 0.002$
Human Enhancer Ensembl $0.744$±$\pm 0.122$ $0.849$±$\pm 0.006$ $0.862$±$\pm 0.008$ $0.893$±$\pm 0.008$ $0.900$±$\pm 0.006$ $0.899$±$\pm 0.003$ $\bf 0.900$±$\pm 0.005$
Human Regulatory $0.872$±$\pm 0.005$ $0.869$±$\pm 0.012$ $0.814$±$\pm 0.211$ $0.872$±$\pm 0.011$ $0.873$±$\pm 0.007$ $\bf 0.877$±$\pm 0.002$ $0.870$±$\pm 0.006$
Human OCR Ensembl $0.698$±$\pm 0.013$ $0.783$±$\pm 0.007$ $0.815$±$\pm 0.002$ $\bf 0.828$±$\pm 0.006$ $0.818$±$\pm 0.006$ $0.825$±$\pm 0.002$ $0.825$±$\pm 0.003$
Human NonTATA Promoters $0.861$±$\pm 0.009$ $0.944$±$\pm 0.002$ $0.933$±$\pm 0.007$ $0.946$±$\pm 0.007$ $0.945$±$\pm 0.010$ $\bf 0.949$±$\pm 0.006$ $0.945$±$\pm 0.006$
@@
Enhancer $0.451$±$\pm 0.108$ $0.516$±$\pm 0.098$ $0.548$±$\pm 0.144$ $0.517$±$\pm 0.117$ $0.546$±$\pm 0.073$ $\bf 0.564$±$\pm 0.022$ $0.535$±$\pm 0.036$
Enhancer types $0.309$±$\pm 0.134$ $0.423$±$\pm 0.051$ $0.424$±$\pm 0.132$ $0.386$±$\pm 0.185$ $0.439$±$\pm 0.054$ $0.462$±$\pm 0.049$ $\bf 0.470$±$\pm 0.025$
Promoter: All $0.954$±$\pm 0.006$ $0.971$±$\pm 0.006$ $\bf 0.976$±$\pm 0.006$ $0.960$±$\pm 0.005$ $0.970$±$\pm 0.004$ $0.969$±$\pm 0.002$ $0.971$±$\pm 0.002$
     NonTATA $0.955$±$\pm 0.010$ $0.972$±$\pm 0.005$ $\bf 0.976$±$\pm 0.005$ $0.959$±$\pm 0.008$ $0.969$±$\pm 0.011$ $0.971$±$\pm 0.003$ $0.971$±$\pm 0.002$
     TATA $0.960$±$\pm 0.023$ $0.955$±$\pm 0.021$ $\bf 0.966$±$\pm 0.013$ $0.944$±$\pm 0.040$ $0.953$±$\pm 0.016$ $0.956$±$\pm 0.010$ $0.958$±$\pm 0.008$
@@ 4.2.3 DNA Long Range Benchmark
Artery Tibial $0.741$ $0.690$ $\bf 0.803$
Adipose Subcutaneous $0.736$ $\bf 0.759$ $0.741$
Cells Cultured fibroblasts $0.639$ $0.690$ $\bf 0.771$
Muscle Skeletal $0.621$ $0.789$ $\bf 0.803$
Nerve Tibial $0.683$ $0.842$ $\bf 0.877$
Skin Not Sun Exposed Suprapubic $0.710$ $0.812$ $\bf 0.875$
Skin Sun Exposed Lower leg $0.700$ $0.692$ $\bf 0.706$
Thyroid $0.612$ $0.703$ $\bf 0.752$
Whole Blood $0.689$ $0.769$ $\bf 0.794$
To further assess our model’s ability to capture long-range dependencies in DNA sequences, we evaluate it on the expression Quantitative Trait Loci (eQTL) prediction task from DNALONGBENCH cheng2025dnalongbench with the sequence length of 450,000. The eQTL task measures whether a nucleotide variant can influence the expression of a target gene based on the sequences of the gene and its surrounding regions.
Due to limited computational resources, we compare our model against the state-of-the-art DNA language model, Caduceus-PH caduceus , which is also trained with the same data scale for fair comparison, and the expert model for this task, Enformer enformer . The results for Enformer are taken directly from DNALONGBENCH cheng2025dnalongbench . For Caduceus-PH, we fine-tune the official HuggingFace-released weights, which are pretrained on sequences of length 131k. Our model is pretrained and fine-tuned using the same setup and sequence length as Caduceus-PH to ensure a fair comparison.
@@ Appendix C Ablation Experiments
To determine the optimal hybrid architecture for DNA sequence modeling that effectively balances local and global attention, we perform ablation experiments on various configurations of the unidirectional encoder (i.e., modules preceding the bidirectional fusion layer). Referring to [21], we also explore the value of the additional mid-attention layer. Specifically, we evaluate the following configurations: 1) mamba and FFN blocks only, 2) mamba and FFN blocks with mid-attention, 3) mamba and FFN blocks with MoE, and 4) mamba and FFN blocks with both mid-attention and MoE. The ratio of MoE to replace FFNs is set to 0.5. In models with mid-attention, the two independent bidirectional Mamba blocks at the 4th layer are replaced with two independent causal Attention blocks implemented using FlashAttention2 [40].
Table 4: Hyperparameter settings for $\sf JanusDNA$ ablation experiments.
JanusDNA (all with Mamba)
ffn midattn+ffn moe+ffn midattn+moe+ffn
Layers 8 8 8 8
Width 148 148 128 128
activated Params (M) 5.973 5.973 6.084 6.080
total Params (M) 5.973 5.973 28.104 28.100
global steps 10k 10k 10k 10k
Expert number of MoE 0 0 16 16
head number of Attention 4 4 4 4
multiple number of FFN width 4 4 4 4
Optimizer AdamW
Optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.95
Learning rate $8\mathrm{e}^{-3}$
LR Scheduler Cosine decay
Weight decay (model) 0.1
The models are pre-trained on sequences of lengths 1024 and 131072, with batch sizes of 128 and 1, respectively, using a single GPU. All other hyperparameters and training settings are consistent with the pre-training setup described earlier. To ensure a consistent number of activated parameters across different models, we adjust the model configurations as Table 4. The training perplexity results for these configurations are shown in Figure 5.
@@
6, indicate that models with mid-attention perform better at a hidden dimension of 32. However, as the hidden dimension increases, models without mid-attention outperform those with mid-attention at a dimension of 72. At a dimension of 128, the performance of models with and without mid-attention becomes comparable. These findings suggest that mid-attention provides diminishing benefits as the hidden dimension grows larger, eventually becoming negligible at higher dimensions.